Q1. Suppose a 3 year bond with a 6% coupon rate that was purchased for $760 and had a promised yield of 8%. Suppose that interest rates increased and the price of the bond declined. Displeased, you sold the bond for 798.8 after having owned it for 1 year. What should be the realized yield ?

Q2. Ahmed has been offered a 10-year bond issued by Homer, Inc., at a price of $800. The bond has a coupon rate of 7 percent and pays the coupon semiannually. Similar bonds in the market will yield 10 percent today.

  1. What should be the price of this bond?
  2. Should Ahmed buy the bonds at the offered price?

* no pics or hand writing

** answers must be acurate and not copied or in excel , must be clarified
Answer #1

1]

realized yield = (sale price - purchase price + coupon payment) / purchase price

coupon payment = par value * coupon rate = $1,000 * 6% = $60

realized yield = ($798.80 - $760 + $60) / $760

realized yield = 13.00%

2]

a]

Price of a bond is the present value of its cash flows. The cash flows are the coupon payments and the face value receivable on maturity

Price of bond is calculated using PV function in Excel :

rate = 10%/2 (Semiannual YTM of bonds = annual YTM / 2)

nper = 10 * 2 (10 years remaining until maturity with 2 semiannual coupon payments each year)

pmt = 1000 * 7% / 2 (semiannual coupon payment = face value * coupon rate / 2)

fv = 1000 (face value receivable on maturity)

PV is calculated to be $813.07

A1 X for =PV(10%/2,10*2,1000*7%/2,1000) E F B C D 1 A ($813.07)

b]

Yes, Ahmed should buy the bonds as they are offered at a price below their fair price